The mile high city of Denver, Colorado is at an altitude of 1600 m above sea-level.

(a) Make the assumption that the density of air (1.2 kg/m3 ) is constant versus height to estimate the air pressure in Denver.

(b) You decide to impress your Rocky Mountain friends by showing how you can use a long straw. How low does the pressure need to be in your lungs to use a straw that is 1.5 m long (assume you are drinking a liquid with density of 1200 kg/m3 )?

(c) Someone says that you would be better off blowing across the top of the straw! If the density of air is 1.0 kg/m3 what velocity of air would be required to acheive a 5.0 kPa reduction in pressure.?

(d) Use the modified form of the Clausius-Clapeyron equation along with a latent heat of vaporization of L = 40.8 kJ/mol to estimate the boiling temperature of water in Denver. The modified equation is ln (p2 /p1) = − L R ( (1 /T2) − (1/T1)) (1) R is the ideal gas constant.
Answer #1

Pressure of atmosphere in mile high city = 0.76 m x 13593 Kg/m^3 x 9.8 m/s^2 - 1600 m x 1.2Kg/m^3 x 9.8 m/s^2

= 101241 - 18816 = 82425 Pascals

Pressure due to liquid column = 1.5m x 1200 Kg/m^3 X 9.8 m/s^2 = 17640 Pascals

Pressure in lungs = 82245 -17640 = 64605 Pascals